# Diffusion approximation of the stationary distribution of a two-level single server queue

### v7

@@ 6.1 Process limit
$p_{2,\ell}=\gamma_{12}p_{2,\ell+1}=\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}},\quad\ell\in[1,\ell_{0}-1].$ (A.11)
Thus, $p_{2,\ell}$ for $\ell\in[0,\ell_{0}-1]$ is determined by $p_{2,\ell_{0}}$. Furthermore, (A.11) can be written by (A.4) as
$p_{2,\ell}=\gamma_{12}^{-1}p_{2,\ell-1}=\gamma_{12}^{-\ell+1}p_{2,1}=\gamma_{12}^{-\ell+1}\rho_{21}p_{2,0},\quad\ell\in[1,\ell_{0}-1].$ (A.12)
Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ at the boundaries states of $S_{j}$ for $j=1,2,3$. We call them boundary probabilities. There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{0}}$ for large $\ell_{0}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}-1}=\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$ by (A.12) and (A.11), so we do not consider it. We first note that, from (A.6),
$p_{2,\ell_{0}}=\gamma_{22}p_{3,\ell_{0}+1}.$ (A.13)
Since $c_{2}\mu p_{\ell_{0}+2}=\lambda_{2}p_{3,\ell_{0}+1}$ by (A.8) for $\ell=\ell_{0}+1$, (A.7) can be written as
$p_{3,\ell_{0}+1}=\rho_{12}p_{1,\ell_{0}}+\rho_{22}p_{2,\ell_{0}},$ (A.14)
and, substituting $p_{3,\ell_{0}+1}=\gamma_{22}^{-1}p_{2,\ell_{0}}$ of (A.13) to this equation, we have
$(\gamma_{22}^{-1}-\rho_{22})p_{2,\ell_{0}}=\rho_{12}p_{1,\ell_{0}}.$
@@
$p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}.$ (A.15)
Thus, the boundary probabilities $p_{1,\ell_{0}}$, $p_{2,\ell_{0}}$ and $p_{3,\ell_{0}+1}$ are explicitly related by (A.13) and (A.15). We also like to see such a relation between $p_{1,0}$ and $p_{1,\ell_{0}}$, but this is not easy.
To resolve this problem, we consider $p_{1,\ell}$ for $\ell\in[0,\ell_{0}-1]$. From (A.2),
$\lambda_{1}p_{1,\ell}-c_{1}\mu p_{1,\ell+1}=\lambda_{1}p_{1,\ell-1}-c_{1}\mu p_{1,\ell}+\lambda_{2}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1],$
@@
$\rho_{11}p_{1,\ell}-p_{1,\ell+1}=\rho_{11}p_{1,\ell-1}-p_{1,\ell}+\rho_{21}p_{2,\ell-1},\qquad\ell\in[1,\ell_{0}-1].$
Summing up both sides of this equation for $\ell\in[1,m]$ for $m\in[1,\ell_{0}-1]$, we have
$\rho_{11}p_{1,m}-p_{1,m+1}=\rho_{21}\sum_{\ell=1}^{m}p_{2,\ell-1},\qquad m\in[0,\ell_{0}-1],$ (A.16)
@@
$\rho_{11}p_{1,m}-p_{1,m+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{\ell=1}^{m-1}\gamma_{12}^{\ell_{0}-\ell}p_{2,\ell_{0}}\right)$
$=\left(1+\rho_{21}\sum_{\ell=1}^{m-1}\gamma_{12}^{-(\ell-1)}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}$
$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[1,\ell_{0}-1].$
@@
$\quad=\rho_{11}^{-(m+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(m\geq 1)\right)\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}},\qquad m\in[0,\ell_{0}-1].$
Summing up both sides of this equation from $m=0$ to $m=k-1\leq\ell_{0}-1$, we have
$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{m=0}^{k-1}\rho_{11}^{-(m+1)}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\sum_{m=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(m-1)}-\rho_{11}^{-(m-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}-\frac{\rho_{11}^{-(k-1)}-1}{\rho_{11}^{-1}-1}\right)\frac{\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}}$
$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}\rho_{11}^{-2}-\frac{\rho_{11}^{-k}-\rho_{11}^{-1}}{1-\rho_{11}}\right)\frac{\rho_{21}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}}{1-\gamma_{12}},\quad k\in[1,\ell_{0}].$
Hence, multiplying both sides of this equation by $\rho_{11}^{k}$, we have
$p_{1,k}$ $=\rho_{11}^{k}p_{1,0}-\frac{1-\rho_{11}^{k}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}p_{2,\ell_{0}}+\frac{1-\rho_{11}^{k-1}}{1-\rho_{11}}\frac{\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}}$
$\quad-\frac{\gamma_{12}^{-(k-1)}-\rho_{11}^{k-1}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}p_{2,\ell_{0}},\qquad k\in[1,\ell_{0}].$ (A.17)
In particular, for $k=\ell_{0}$, substituting $p_{2,\ell_{0}}=\rho_{11}p_{1,\ell_{0}}$ of (A.15), we have
$\rho_{11}^{\ell_{0}}\frac{p_{1,0}}{p_{1,\ell_{0}}}$ $=1+\frac{1-\rho_{11}^{\ell_{0}}}{1-\rho_{11}}\gamma_{12}^{\ell_{0}-1}\rho_{11}-\frac{1-\rho_{11}^{\ell_{0}-1}}{1-\rho_{11}}\frac{\rho_{11}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{0}}$
$\quad+\frac{\rho_{11}\gamma_{12}-\rho_{11}^{\ell_{0}}\gamma_{12}^{\ell_{0}}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}.$ (A.18)
Thus, $p_{1,0}$ is explicitly related to $p_{1,\ell_{0}}$.
@@
$g^{(r)}_{j}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}}e^{r\theta k}p^{(r)}_{j,k},\quad\theta\in\mathbb{R},j=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{0}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
@@
$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{0}r\theta}\sum_{\ell=1}^{\ell^{(r)}_{0}}e^{-(\ell^{(r)}_{0}-\ell)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-\ell}p^{(r)}_{2,\ell^{(r)}_{0}}$
$=\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}+e^{r\theta}\frac{e^{\ell^{(r)}_{0}r\theta}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{0}},$ (A.19)
$g^{(r)}_{3}(\theta)$ $=\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{0}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{0}+1}.$ (A.20)
On the other hand, computing $g^{(r)}_{1}(\theta)$ is a bit complicated. To ease it, we introduce the following functions.
$f^{(r)}_{1}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}e^{r\theta k},\qquad f^{(r)}_{\rho_{1}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}\rho^{(r)}_{1})^{k},\qquad f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{0}-1}(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k},$
then
$f^{(r)}_{1}(\theta)=\frac{1-(e^{r\theta})^{\ell^{(r)}_{0}}}{1-e^{r\theta}}1(\theta\not=0)+\ell^{(r)}_{0}1(\theta=0),\qquad f^{(r)}_{\rho_{1}}(\theta)=\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}}{1-e^{r\theta}\rho^{(r)}_{1}},$
$f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\frac{e^{r\theta\ell^{(r)}_{0}}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}}{e^{r\theta}-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{-\ell^{(r)}_{0}+1}.$
Using these function, we compute $g^{(r)}_{1}(\theta)$ from (A.2) as given below.
$g^{(r)}_{1}(\theta)$ $=\sum_{k=0}^{\ell^{(r)}_{0}}(e^{r\theta}\rho^{(r)}_{1})^{k}p^{(r)}_{1,0}-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta k}-(e^{r\theta}\rho^{(r)}_{1})^{k}}{1-\rho^{(r)}_{1}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad+\sum_{k=1}^{\ell^{(r)}_{0}}\frac{e^{r\theta(k-1)}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{1-\rho^{(r)}_{1}}\frac{e^{r\theta}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad-\sum_{k=1}^{\ell^{(r)}_{0}}\frac{(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k-1}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell_{0}}$
$=\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{0}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(\theta)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad+\left(f^{(r)}_{1}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$
$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}p^{(r)}_{2,\ell^{(r)}_{0}}$ (A.21)
@@
$(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}=O(\gamma_{12}^{\ell_{0}/r}),\qquad\mbox{as }r\downarrow 0.$ (A.22)
Hence, $(\gamma^{(r)}_{12})^{\ell^{(r)}_{0}}$ decays exponentially fast as the function of $1/r$. On the other hand,
@@
$\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{\ell^{(r)}_{0}}=\lim_{r\downarrow 0}\left(1-\beta_{1}r+o(r)\right)^{\ell_{0}/r+o(r)}=e^{-\beta_{1}\ell_{0}},$ (A.24)
@@
$\lim_{r\downarrow 0}\frac{p^{(r)}_{2,\ell_{0}}}{p^{(r)}_{1,\ell_{0}}}=1,\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{3,\ell_{0}+1}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}},\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{0}}.$ (A.25)
@@
$\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{0}}}$ $=\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{-\ell^{(r)}_{0}}\left(1+\frac{\rho^{(r)}_{1}\gamma^{(r)}_{12}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}\right).$

### v8

@@ 6.1 Process limit
$p_{2,\ell}=\gamma_{12}p_{2,\ell+1}=\gamma_{12}^{\ell_{1}-\ell}p_{2,\ell_{1}},\quad\ell\in[1,\ell_{1}-1].$ (A.11)
Thus, $p_{2,\ell}$ for $\ell\in[0,\ell_{1}-1]$ is determined by $p_{2,\ell_{1}}$. Furthermore, (A.11) can be written by (A.4) as
$p_{2,\ell}=\gamma_{12}^{-1}p_{2,\ell-1}=\gamma_{12}^{-\ell+1}p_{2,1}=\gamma_{12}^{-\ell+1}\rho_{21}p_{2,0},\quad\ell\in[1,\ell_{1}-1].$ (A.12)
Let us consider relations among the stationary probabilities, $p_{1,0},p_{1,\ell_{1}}$, $p_{2,\ell_{1}}$ and $p_{3,\ell_{1}+1}$ at the boundaries states of $S_{j}$ for $j=1,2,3$. We call them boundary probabilities. There is one more boundary probability, $p_{2,0}$, but this probability is ignorable compared with $p_{2,\ell_{1}}$ for large $\ell_{1}$ as $\rho_{21}p_{2,0}=\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}-1}=\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}$ by (A.12) and (A.11), so we do not consider it. We first note that, from (A.6),
$p_{2,\ell_{1}}=\gamma_{22}p_{3,\ell_{1}+1}.$ (A.13)
Since $c_{2}\mu p_{\ell_{1}+2}=\lambda_{2}p_{3,\ell_{1}+1}$ by (A.8) for $\ell=\ell_{1}+1$, (A.7) can be written as
$p_{3,\ell_{1}+1}=\rho_{12}p_{1,\ell_{1}}+\rho_{22}p_{2,\ell_{1}},$ (A.14)
and, substituting $p_{3,\ell_{1}+1}=\gamma_{22}^{-1}p_{2,\ell_{1}}$ of (A.13) to this equation, we have
$(\gamma_{22}^{-1}-\rho_{22})p_{2,\ell_{1}}=\rho_{12}p_{1,\ell_{1}}.$
@@
$p_{2,\ell_{1}}=\rho_{11}p_{1,\ell_{1}}.$ (A.15)
Thus, the boundary probabilities $p_{1,\ell_{1}}$, $p_{2,\ell_{1}}$ and $p_{3,\ell_{1}+1}$ are explicitly related by (A.13) and (A.15). We also like to see such a relation between $p_{1,0}$ and $p_{1,\ell_{1}}$, but this is not easy.
To resolve this problem, we consider $p_{1,\ell}$ for $\ell\in[0,\ell_{1}-1]$. From (A.2),
$\lambda_{1}p_{1,\ell}-c_{1}\mu p_{1,\ell+1}=\lambda_{1}p_{1,\ell-1}-c_{1}\mu p_{1,\ell}+\lambda_{2}p_{2,\ell-1},\qquad\ell\in[1,\ell_{1}-1],$
@@
$\rho_{11}p_{1,\ell}-p_{1,\ell+1}=\rho_{11}p_{1,\ell-1}-p_{1,\ell}+\rho_{21}p_{2,\ell-1},\qquad\ell\in[1,\ell_{1}-1].$
Summing up both sides of this equation for $\ell\in[1,m]$ for $m\in[1,\ell_{1}-1]$, we have
$\rho_{11}p_{1,m}-p_{1,m+1}=\rho_{21}\sum_{\ell=1}^{m}p_{2,\ell-1},\qquad m\in[0,\ell_{1}-1],$ (A.16)
@@
$\rho_{11}p_{1,m}-p_{1,m+1}$ $=\rho_{21}\left(\rho_{21}^{-1}p_{2,1}+\sum_{\ell=1}^{m-1}\gamma_{12}^{\ell_{1}-\ell}p_{2,\ell_{1}}\right)$
$=\left(1+\rho_{21}\sum_{\ell=1}^{m-1}\gamma_{12}^{-(\ell-1)}\right)\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}$
$=\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}\right)\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}},\qquad m\in[1,\ell_{1}-1].$
@@
$\quad=\rho_{11}^{-(m+1)}\left(1+\rho_{21}\frac{(\gamma_{12}^{-(m-1)}-1)\gamma_{12}}{1-\gamma_{12}}1(m\geq 1)\right)\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}},\qquad m\in[0,\ell_{1}-1].$
Summing up both sides of this equation from $m=0$ to $m=k-1\leq\ell_{1}-1$, we have
$p_{1,0}-\rho_{11}^{-k}p_{1,k}=\sum_{m=0}^{k-1}\rho_{11}^{-(m+1)}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\sum_{m=1}^{k-1}\frac{(\rho_{11}\gamma_{12})^{-(m-1)}-\rho_{11}^{-(m-1)}}{1-\gamma_{12}}\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}$
$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}-\frac{\rho_{11}^{-(k-1)}-1}{\rho_{11}^{-1}-1}\right)\frac{\rho_{11}^{-2}\rho_{21}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}}{1-\gamma_{12}}$
$\quad=\frac{\rho_{11}^{-k}-1}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\left(\frac{(\rho_{11}\gamma_{12})^{-(k-1)}-1}{(\rho_{11}\gamma_{12})^{-1}-1}\rho_{11}^{-2}-\frac{\rho_{11}^{-k}-\rho_{11}^{-1}}{1-\rho_{11}}\right)\frac{\rho_{21}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}}{1-\gamma_{12}}$
for $k\in[1,\ell_{1}]$. Hence, multiplying both sides of this equation by $\rho_{11}^{k}$, we have
$p_{1,k}$ $=\rho_{11}^{k}p_{1,0}-\frac{1-\rho_{11}^{k}}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}p_{2,\ell_{1}}+\frac{1-\rho_{11}^{k-1}}{1-\rho_{11}}\frac{\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}}$
$\quad-\frac{\gamma_{12}^{-(k-1)}-\rho_{11}^{k-1}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{1}}p_{2,\ell_{1}},\qquad k\in[1,\ell_{1}].$ (A.17)
In particular, for $k=\ell_{1}$, substituting $p_{2,\ell_{1}}=\rho_{11}p_{1,\ell_{1}}$ of (A.15), we have
$\rho_{11}^{\ell_{1}}\frac{p_{1,0}}{p_{1,\ell_{1}}}$ $=1+\frac{1-\rho_{11}^{\ell_{1}}}{1-\rho_{11}}\gamma_{12}^{\ell_{1}-1}\rho_{11}-\frac{1-\rho_{11}^{\ell_{1}-1}}{1-\rho_{11}}\frac{\rho_{11}\rho_{21}}{1-\gamma_{12}}\gamma_{12}^{\ell_{1}}$
$\quad+\frac{\rho_{11}\gamma_{12}-\rho_{11}^{\ell_{1}}\gamma_{12}^{\ell_{1}}}{(\rho_{11}\gamma_{12})^{-1}-1}\frac{\rho_{11}^{-1}\rho_{21}}{1-\gamma_{12}}.$ (A.18)
Thus, $p_{1,0}$ is explicitly related to $p_{1,\ell_{1}}$.
@@
$g^{(r)}_{j}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}}e^{r\theta k}p^{(r)}_{j,k},\quad\theta\in\mathbb{R},j=1,2,\qquad g^{(r)}_{3}(\theta)=\sum_{k=\ell^{(r)}_{1}+1}^{\infty}e^{r\theta k}p^{(r)}_{3,k},\quad\theta\leq 0,$
@@
$g_{2}^{(r)}(\theta)$ $=p^{(r)}_{2,0}+e^{\ell^{(r)}_{1}r\theta}\sum_{\ell=1}^{\ell^{(r)}_{1}}e^{-(\ell^{(r)}_{1}-\ell)r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-\ell}p^{(r)}_{2,\ell^{(r)}_{1}}$
$=\left((\rho^{(r)}_{21})^{-1}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}+e^{r\theta}\frac{e^{\ell^{(r)}_{1}r\theta}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{e^{r\theta}-\gamma^{(r)}_{12}}\right)p^{(r)}_{2,\ell^{(r)}_{1}},$ (A.19)
$g^{(r)}_{3}(\theta)$ $=\frac{1}{1-e^{r\theta}\rho^{(r)}_{2}}e^{(\ell^{(r)}_{1}+1)r\theta}p^{(r)}_{3,\ell^{(r)}_{1}+1}.$ (A.20)
On the other hand, we compute $g^{(r)}_{1}(\theta)$ from (A.3) as given below.
$g^{(r)}_{1}(\theta)$ $=\sum_{k=0}^{\ell^{(r)}_{1}}(e^{r\theta}\rho^{(r)}_{1})^{k}p^{(r)}_{1,0}-\sum_{k=1}^{\ell^{(r)}_{1}}\frac{e^{r\theta k}-(e^{r\theta}\rho^{(r)}_{1})^{k}}{1-\rho^{(r)}_{1}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}p^{(r)}_{2,\ell^{(r)}_{1}}$
$\quad+\sum_{k=1}^{\ell^{(r)}_{1}}\frac{e^{r\theta(k-1)}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{1-\rho^{(r)}_{1}}\frac{e^{r\theta}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
$\quad-\sum_{k=1}^{\ell^{(r)}_{1}}\frac{(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k-1}-(e^{r\theta}\rho^{(r)}_{1})^{k-1}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell_{1}}.$ (A.21)
To ease further compuations, we introduce the following functions.
$f^{(r)}_{1}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}-1}e^{r\theta k},\qquad f^{(r)}_{\rho_{1}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}-1}(e^{r\theta}\rho^{(r)}_{1})^{k},\qquad f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\sum_{k=0}^{\ell^{(r)}_{1}-1}(e^{r\theta}(\gamma^{(r)}_{12})^{-1})^{k},$
which are computed as
$f^{(r)}_{1}(\theta)=\frac{1-(e^{r\theta})^{\ell^{(r)}_{1}}}{1-e^{r\theta}}1(\theta\not=0)+\ell^{(r)}_{1}1(\theta=0),\qquad f^{(r)}_{\rho_{1}}(\theta)=\frac{1-(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}}}{1-e^{r\theta}\rho^{(r)}_{1}},$
$f^{(r)}_{\gamma^{-1}_{12}}(\theta)=\frac{e^{r\theta\ell^{(r)}_{1}}-(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}}{e^{r\theta}-\gamma^{(r)}_{12}}(\gamma^{(r)}_{12})^{-\ell^{(r)}_{1}+1}.$
Then, using these function, (A.3) becomes
$g^{(r)}_{1}(\theta)$ $=\left(f^{(r)}_{\rho_{1}}(\theta)+(e^{r\theta}\rho^{(r)}_{1})^{\ell^{(r)}_{1}}\right)p^{(r)}_{1,0}-\left(f^{(r)}_{1}(\theta)-\rho^{(r)}_{1}f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}-1}}{1-\rho^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
$\quad+\left(f^{(r)}_{1}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}\rho^{(r)}_{21}}{(1-\rho^{(r)}_{1})(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$
$\quad-\left(f^{(r)}_{\gamma^{-1}_{12}}(\theta)-f^{(r)}_{\rho_{1}}(\theta)\right)\frac{e^{r\theta}(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{((\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1)(1-\gamma^{(r)}_{12})}(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}p^{(r)}_{2,\ell^{(r)}_{1}}$ (A.22)
@@
$(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}=O(\gamma_{12}^{\ell_{1}/r}),\qquad\mbox{as }r\downarrow 0.$ (A.23)
Hence, $(\gamma^{(r)}_{12})^{\ell^{(r)}_{1}}$ decays exponentially fast as the function of $1/r$. On the other hand,
@@
$\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{\ell^{(r)}_{1}}=\lim_{r\downarrow 0}\left(1-\beta_{1}r+o(r)\right)^{\ell_{1}/r+o(r)}=e^{-\beta_{1}\ell_{1}},$ (A.25)
@@
$\lim_{r\downarrow 0}\frac{p^{(r)}_{2,\ell_{1}}}{p^{(r)}_{1,\ell_{1}}}=1,\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{3,\ell_{1}+1}}{p^{(r)}_{1,\ell_{1}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}},\qquad\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{1}}}=\frac{\lambda_{1}+\lambda_{2}}{\lambda_{2}}e^{\beta_{1}\ell_{1}}.$ (A.26)
@@
$\lim_{r\downarrow 0}\frac{p^{(r)}_{1,0}}{p^{(r)}_{1,\ell_{1}}}$ $=\lim_{r\downarrow 0}(\rho^{(r)}_{1})^{-\ell^{(r)}_{1}}\left(1+\frac{\rho^{(r)}_{1}\gamma^{(r)}_{12}}{(\rho^{(r)}_{1}\gamma^{(r)}_{12})^{-1}-1}\frac{(\rho^{(r)}_{1})^{-1}\rho^{(r)}_{21}}{1-\gamma^{(r)}_{12}}\right).$